OpenStudy (anonymous):

Solve the proportion: x/2 = 5/15 x =

OpenStudy (anonymous):

is x 6

OpenStudy (anonymous):

well first you can simplify 5/15 to 1/3

OpenStudy (anonymous):

then you set x/2=1/3 and just simply cross multiply

OpenStudy (anonymous):

i got x=2/3 o.o

OpenStudy (anonymous):